OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

one second let me type the equation.

OpenStudy (alexandervonhumboldt2):

ooh no so many helpers i want to help

OpenStudy (alexandervonhumboldt2):

i viewing your question

OpenStudy (anonymous):

\[if f(x)=2x/2x+1 and g(x)=1/2x+1, find f(x)+g(x)\]

OpenStudy (alexandervonhumboldt2):

2x/2x+1+1/2x+1?

OpenStudy (alexandervonhumboldt2):

1+1+1/2x+1=3+1/2x then

OpenStudy (anonymous):

yea